GXY Corp. has a WACC of 15%, a cost of debt capital of 5%, and a market value debt-to-equity ratio of 0.10. GXY Corp. is not subject to taxation. Suppose that GXY Corp. increased it market value debt-to-equity ratio to 0.35, What will be the change in GXY Corp’s WACC? Enter your answer as a percent; do not include the % sign. Round your final answer to two decimals.
Answer #1


Solution: WACC=rd*(1-t)*wd+rs*ws 0.15=0.05*(1-0)*(0.1/(0.1+1))+rs* (1/(0.1+1)) 0.15-0.05*(0.1/1.1)+rs *(1/1.1) rs=(0.15-0.05*